OpenStudy (anonymous):

the quantity of 2 c to the 4th power times the quantity of negative of c squared times the quantity of 4 c to the ninth power =

OpenStudy (anonymous):

\[2c^4\times (-c)^2\times 4c^9\] or \[2c^4\times -c^2\times 4c^9\]

OpenStudy (anonymous):

\[(2c ^{4}) x (-c ^{2}) x (4c ^{9})\]

OpenStudy (anonymous):

add all the exponents

OpenStudy (anonymous):

and put a \(-8\) out front since \(2\times (-1)\times 4=-8\)

OpenStudy (anonymous):

ok then what do i do

OpenStudy (anonymous):

add the exponents \(4+2+9=?\)

OpenStudy (anonymous):

15?